1. Use MM propositions I and II to solve a - c. Show your work. Your firm is currently financed with 50% debt and 50% equity. You have been told that the cost of unlevered equity is 15% and the cost of debt at D/E ratios ranging from 0 to 3.0 is 5%.  
  1. Calculate the cost of levered equity.

  1. Calculate the WACC

c. What would the WACC be if your firm was financed with 75% debt and 25% equity?
Answer #1

Answer:

Given

Debt D=50%

Equity E=50%

D/E=1

Cost of Debt Rd =5%

Cost unlevered equity Ru=15%

1)

Cost of Levered equity Re=Ru+D/E*(Ru-Rd)=

Re=15%+1*(15%-5%)=25%

2)

WACC=D*Rd+E*Re=50%*5%+50%*25%=15%

3)

Debt D=75%

Equity E=25%

D/E=3

Cost of Debt Rd =5%

Cost unlevered equity Ru=15%

Cost of Levered equity Re=Ru+D/E*(Ru-Rd)=

Re=15%+3*(15%-5%)=45%

WACC=D*Rd+E*Re=75%*5%+25%*45%=15%